# On The Heine-Borel Property and Minimum Enclosing Balls

### v1

@@ Abstract
In this paper, we contribute a proof that minimum radius balls over metric spaces with the Heine-Borel property are always LP type. Additionally, we prove that weak metric spaces, those without symmetry, also have this property if we fix the direction in which we take their distances from the centers of the balls. We use this to prove that the minimum radius ball problem is LP type in the Hilbert and Thompson metrics and Funk weak metric. In doing so, we contribute a proof that the topology induced by the Thompson metric coincides with the Hilbert. We provide explicit primitives for computing the minimum radius ball in the Hilbert metric.
@@ 1 Introduction
The concept of an LP-type problem originates with the work of Micah Sharir and Emo Welzl in 1992 in their paper “A combinatorial bound for linear programming and related problems” [25]. Since then, a variety of problems have been shown to be LP type, including minimum radius balls in the Euclidean metric [16] and spaces with Bregman divergences [17], finding the closest distance between two convex polygons [16], various game-theoretic games [11] such as simple stochastic games and parity games. The analysis of LP-type problems has been a continuous area of study in computational geometry since their inception. We contribute to this by proving that determining the minimum radius ball in any metric space satisfying the Heine-Borel property is LP-type. We provide an example with the Hilbert and Thompson metrics. We also show that this holds for weak metric spaces, that is, where the distance function fails to be symmetric.
@@
The Hilbert metric originates from the work of David Hilbert in 1895 in relation to Hilbert’s fourth problem [12]. It presents a non-Euclidean metric in which the triangle inequality is not strict. The Hilbert metric generalizes the Cayley-Klein model of hyperbolic geometry to arbitrary convex bodies in a $d$-dimensional space. Definitions will be given in Section 2. Given a convex body, $\Omega\subset\mathbb{R}^{d}$, the Hilbert metric has many desirable properties, such as the fact that straight lines are geodesics and are preserved under projective transformations. In the probability simplex, it provides a natural distance between discrete probability distributions, as shown by the works of Nielsen and Sun [19, 20]. For an excellent resource on Hilbert geometry, see the “Handbook of Hilbert Geometry” [22], by Papadopoulos and Troyanov.
@@
The Hilbert metric has seen recent use in a diverse set of fields, especially that of convexity approximation. This is in particular due to its relation with Macbeath regions (which are equivalent to Hilbert balls up to a scaling factor) [1, 2] and the flag approximability of polytopes [28]. Other fields in which the Hilbert metric has been used are quantum information theory on convex cones defined by various operators [23], machine learning in the form of clustering and graph embeddings [19, 20], optimal mass transport [5], and a variety of situations in real analysis [22]. Due to its many uses, various algorithms from classical computational geometry have been modified for use in the Hilbert metric, including Voronoi diagrams [10, 4] and Delaunay triangulations [9]. We expand on these works by contributing an algorithm for minimum enclosing balls in the Hilbert polygonal geometry.
@@
The Thompson metric was defined by A. C. Thompson in 1963 as an alternative to the Hilbert metric for its applications in analysis[26]. Like the Hilbert, it provides a metric space over convex bodies and has a similar geometry. Its primary uses are in analysis, in particular as a metric on cones [6, 13, 14, 15]. We contribute a proof that minimum radius balls in this metric are LP type. Additionally, we contribute the fact that Thompson balls are, like Macbeath Regions, also equivalent to Hilbert balls up to a scaling factor, and therefore induce the same topology on convex bodies.
@@
The Funk weak metric was defined by Paul Funk in 1929 [8]. The Funk metric is a weak metric space that can be used to define both the Hilbert and the Thompson metrics, and is often studied in their context [21, 27] and in the context of flags of polytopes [7]. Because the Funk metric is non-symmetric, it induces a reverse metric called the reverse Funk metric. We contribute a proof showing that minimum radius balls in these weak metrics are LP type.
@@ 2.1 Metric Spaces
When all the above properties hold except symmetry, the space is called a weak metric. Given a metric space, $(\mathbf{X},d)$, the closed ball of radius $r$ around a point $p$ generalizes the Euclidean idea of a circle to an arbitrary metric space. We define it here:
@@
$B(p,r)\leavevmode\nobreak\ =\leavevmode\nobreak\ \{q\in\mathbf{X}\mid d(p,q)\leq r\}.$
@@ 2.2 The Hilbert and Thompson Metrics, and the Funk Weak Metric
The Hilbert and Thompson metrics, and the Funk weak metrics, are defined on the points in the interior of a convex body $\Omega$ in $\mathbb{R}^{d}$, where $\Omega$ is a closed and bounded convex set. Unless otherwise stated, we assume that $\Omega$ is a convex polygon with $m$ sides. We let $\partial\kern 1.0pt\Omega$ refer to the boundary of $\Omega$. Given any pair of points $p$ and $q$ within the interior of $\Omega$, we define $\overline{pq}$ to be the cord of $\Omega$ through those two points, and $\chi(p,q)$ to be the directed ray from $p$ through $q$. Unless otherwise stated, when we take $p,q\in\Omega$, we mean the interior of $\Omega$.
@@
Given two points $p,q$ in a convex polygon $\Omega$ in $\mathbb{R}^{d}$ such that $\chi(p,q)$ intersects $\partial\kern 1.0pt\Omega$ at a point $q^{\prime}$ on $\partial\kern 1.0pt\Omega$, we define the Funk weak metric to be:
@@
Given two distinct points $p,q$ in a convex polygon $\Omega$ in $\mathbb{R}^{d}$, let $p^{\prime}$ and $q^{\prime}$ denote the endpoints of the chord $\overline{pq}$ on $\partial\kern 1.0pt(\Omega)$, so that the points lie in order $\langle p^{\prime},p,q,q^{\prime}\rangle$, the Hilbert distance between $p$ and $q$, $H_{\Omega}(p,q)$, is:
@@
Since the product in the definition of the Hilbert metric is the cross ratio, which is preserved under projective transformations, it follows that the Hilbert metric is invariant under projective transformations [18]. It is also worth noting that straight lines are geodesics in the Hilbert metric, though not all geodesics are straight lines, and that the Hilbert metric satisfies all the properties of a metric, including symmetry and the triangle inequality. Nielsen and Shao showed how to algorithmically compute Hilbert balls with the aid of spokes [18]. Given $p\in\Omega$ and a vertex $v$ of $\Omega$, a spoke through $p$ from $v$ is $\overline{pv}$.
Succinctly, a Hilbert Ball of radius $r$ around a point $p$ can be constructed by computing all the points that are a distance of $r$ away from $p$ along the spokes of $p$ from all the vertices of $\Omega$ and then forming their convex hull (see Figure 1(c)).
@@
Balls in the Hilbert geometry of an $m$-sided convex polygon $\Omega$ have between $m$ and $2m$ sides.
@@
Given a bounded closed convex body $\Omega$ in $\mathbb{R}^{d}$ and two points $p,q\in\operatorname{int}(\Omega)$, let $p^{\prime}$ and $q^{\prime}$ denote the endpoints of the chord $\overline{pq}$ on $\partial\kern 1.0pt(\Omega)$, so that the points lie in order $\langle p^{\prime},p,q,q^{\prime}\rangle$, the Thompson distance between $p$ and $q$ $T_{\Omega}(p,q)$ is:
$T_{\Omega}(p,q)\leavevmode\nobreak\ =\leavevmode\nobreak\ \max\left(\ln\frac{\|q-p^{\prime}\|}{\|p-p^{\prime}\|},\ln\frac{\|p-q^{\prime}\|}{\|q-q^{\prime}\|}\right),$
@@
Balls in the Thompson geometry of an $m$-sided convex polygon $\Omega$ are convex polygons with between $m$ and $2m$ sides.
Proof 2.8.
It is evident from its definition that the Thompson ball of radius $r$ centered at a point $x$ is the intersection of the Funk and reverse-Funk balls of radius $r$ centered at $x$. Since the Funk and reverse-Funk balls are homothets of $\Omega$ and $-\Omega$, respectively [21] (see Figure 1(a) and (d)), and since $\Omega$ has $m$ sides, it follows that the number of sides of the Thompson ball ranges from $m$ to $2m$ (see Figure 1(b)).
@@ 3 Minimum Enclosing Radius Balls and the Heine-Borel Property
Figure 2: This is $\mathbb{R}^{2}$ with several closed balls removed. There is no Euclidean minimum radius ball through $x$, $y$, and $z$.
@@
For an example of a metric space without the minimum ball property, see Figure 2. In this situation, a bunch of closed disks were removed from $\mathbb{R}^{2}$ with the Euclidean distance.
@@
$S:=\{s\in\mathbb{R}^{\geq 0}\mid\text{ there exists a closed ball of radius }s\text{ enclosing }H\}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ r:=\inf S.$
Note that $S$ is bounded below by 0 and is nonempty because $H$ is bounded as a consequence of being finite. So the infimum is defined and finite. Thus, to show there exists a ball of minimal radius, we want to show that there exists a ball of radius $r$ enclosing $H$. Define the set of centers $C_{n}:=\{p\in\mathbf{X}:H\subset B(p,r+1/n)\}$. For all $n\in\mathbb{N}$ we have $C_{n}\supset C_{n+1}$ because for all $p\in C_{n+1}$ we have
$H\leavevmode\nobreak\ \subset\leavevmode\nobreak\ B\left(p,r+\frac{1}{n+1}\right)\leavevmode\nobreak\ \subset\leavevmode\nobreak\ B\left(p,r+\frac{1}{n}\right),$
and thus, $p\in C_{n}$ as well. Observe that each $C_{n}$ is bounded since for any $h\in H$ and any $p\in C_{n}$, $d(p,h)\leq r+1/n$.
We will now show that each $C_{n}$ is closed. Let $h\in H$ and $\{p_{i}\}_{i\in\mathbb{N}}\subset C_{n}$ such that $p_{i}\rightarrow p$, $p\in\mathbf{X}$. Then for any $\varepsilon>0$, there exists $i$ such that $d(p_{i},p)<\varepsilon$. So by the triangle inequality,
$d(p,h)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ d(p,p_{i})+d(p_{i},h)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \left(r+\frac{1}{n}\right)+\varepsilon,$
for any $\varepsilon>0$, so $d(p,h)\leq r+1/n$. So $h\in B(p,r+1/n)$ for any $h\in H$. Thus $p\in C_{n}$. So $C_{n}$ contains all of its limit points and is thus closed.
Define $D:=\bigcap_{n\in\mathbb{N}}C_{n}$. Since we are assuming the Heine-Borel property and showed each $C_{n}$ is closed and bounded, each $C_{n}$ is compact. Because $\{C_{n}\}$ is a decreasing nested sequence of compact sets of $\mathbf{X}$ with the metric topology, by Cantor’s intersection theorem $D$ is nonempty [24].
@@
Note that the converse does not hold. As a counterexample, any infinite set equipped with the discrete metric does not satisfy the Heine-Borel property, yet it satisfies the minimum ball property. Additionally, note that in the above proof of Proposition 3.3. we did not use the symmetric property of the metric space, only the triangle inequality. As such, given the same definition of a ball Definition 2.2 and Definition 3.1, e.g. fixing the direction of the ball, we contribute the following corollary.
@@
Now that we have done this, we introduce the criteria outlined for being an LP-type problem by Sharir and Welzl [25].
@@
Let $(\mathbf{X},d)$ be a metric space satisfying the minimum ball property. Let $\mathbf{X}$ have some well ordering $\preceq$. Take a finite set $H\subset\mathbf{X}$. We define a function $f:2^{H}\rightarrow\mathbb{R}\times\mathbf{X}$, where,
@@
Here, $f(G)=(\hat{r},\hat{p})$ gives us a unique minimum radius enclosing ball of $G$, by taking the minimum center with respect to the well ordering. We call this ball $B_{G}=B(\hat{p},\hat{r})$. Note that if $f(G)=(r,p)$, there may exist other points $p^{\prime}$ such that $G\subset B(r,p^{\prime})$ as well; however, $f$ defines only one minimum radius enclosing ball. Note also that if $\mathbf{X}=\mathbb{R}^{d}$ then by Proposition 3.10. we can take the lexicographic usual total order and $f$ will still be well-defined.
@@
Finding the minimum radius ball of a set of points in a metric space that satisfies the minimum ball property is LP type.
@@
To see that $f$ satisfies locality, consider a finite set $H\subset\mathbf{X}$. Take $F,G\subset H$ and $x\in H$ such that $F\subset G\subset H$ and $f(F)=f(G)=f(F\cup\{x\})$. It follows that $B_{F}=B_{G}=B_{F\cup\{x\}}$. Hence, $x\in B_{G}$ and $F\subset G\cup\{x\}\subset B_{G}=B_{F}$. Thus, by monotonicity,
$f(F)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ f(G\cup\{x\})\leavevmode\nobreak\ \leq\leavevmode\nobreak\ f(G)\leavevmode\nobreak\ =\leavevmode\nobreak\ f(F),$
and we conclude that $f(F)=f(G\cup\{x\})$.
@@
Finding the minimum radius ball of a set of points in a weak metric space that satisfies the minimum ball property is LP type.
Note that a well ordering is not always necessary and can be replaced with a total ordering, such as a lexicographic ordering.
@@
If $X=\mathbb{R}^{k}$ for $k\in\mathbb{N}$, then the well order $\preceq$ on $\mathbf{X}$ can be replaced with the lexicographical order $\leq$” over coordinates in $\mathbb{R}^{k}$.
@@
Recall, the following notation and results from Proposition 3.3.
$C_{n}:=\{p\in\mathbf{X}:H\subset B(p,r+1/n)\}\qquad\text{and}\qquad D:=\bigcap_{n\in\mathbb{N}}C_{n}.$
We saw that each $C_{n}$ is closed and bounded with respect to the metric topology. Thus, it is compact in the standard (metric) topology in $\mathbb{R}^{k}$, and thus $D$, the set of centers of minimum radius enclosing balls is also nonempty and compact in this topology.

### v2

@@ Abstract
In this paper, we contribute a proof that the problem of determining the minimum radius balls over metric spaces with the Heine-Borel property is always LP-type. Additionally, we prove that weak metric spaces, those without symmetry, also have this property if we fix the direction in which we take their distances from the centers of the balls. We use this to prove that the minimum radius ball problem is also LP-type in the Thompson metrics and Funk weak metric. We finally examine the LP algorithm and explicit primitives for computing the minimum radius ball in the Hilbert metric.
@@ 1 Introduction
The concept of an LP-type problem originates with the work of Micah Sharir and Emo Welzl in 1992 in their paper "A combinatorial bound for linear programming and related problems"[26]. Since then, a variety of problems have been shown to be LP-type, including minimum radius balls in the Euclidean metric [17] and spaces with Bregman divergences [18], finding the closest distance between two convex polygons [17], various game-theoretic games [11] such as simple stochastic games and parity games. The analysis of LP-type problems has been a continuous area of study in computational geometry since their inception. We contribute to this by proving that determining the minimum radius ball in any metric space satisfying the Heine-Borel property is LP-type. We provide an example with the Hilbert and Thompson metrics. We also show that this holds for weak metric spaces, that is, where the distance function fails to be symmetric.
@@
The Hilbert metric originates from the work of David Hilbert in 1895 in relation to Hilbert’s fourth problem [12]. It presents a non-Euclidean metric in which the triangle inequality is not strict. The Hilbert metric generalizes the Cayley-Klein model of hyperbolic geometry to arbitrary convex bodies in an $n$-dimensional space. Definitions will be given in Section 2. Given a convex body, $\Omega\subset\mathbb{R}^{n}$, the Hilbert metric has many desirable properties, such as the fact that straight lines are geodesics and are preserved under projective transformations. In the probability simplex, it provides a natural distance between discrete probability distributions, as shown by the works of Nielsen and Sun [20, 21]. For an excellent resource on Hilbert geometry, see "From Funk to Hilbert Geometry" or the "Handbook of Hilbert Geometry", both by Papadopoulos and Troyanov[23, 22].
@@
The Hilbert metric has seen recent use in a diverse set of fields, especially in that of convexity approximation. This is in particular due to its relationship with Macbeath regions (which are equivalent to Hilbert balls up to a scaling factor) [1, 2] and the flag approximability of polytopes [29]. Other fields in which the Hilbert metric has been used are quantum information theory on convex cones defined by various operators[24], machine learning in the form of clustering and graph embeddings [20, 21], optimal mass transport [5], and a variety of situations in real analysis [13]. Due to its many uses, various algorithms from classical computational geometry have been modified for use in the Hilbert metric, including Voronoi diagrams [10, 4] and Delaunay triangulations [9]. We expand on these works by contributing an algorithm for minimum enclosing balls in the Hilbert polygonal geometry.
@@
The Thompson metric was defined by A. C. Thompson in 1963 as an alternative to the Hilbert metric for its applications in analysis[27]. Like the Hilbert, it provides a metric space over convex bodies and has a similar geometry. Its primary uses are in analysis, in particular as a metric on cones [6, 14, 15, 16]. We contribute a proof that minimum radius balls in this metric are LP-type. Additionally, we contribute the fact that Thompson balls are, like Macbeath Regions, also equivalent to Hilbert balls up to a scaling factor, and therefore induce the same topology on convex bodies.
@@
The Funk weak metric was defined by Paul Funk in 1929 [8]. The Funk metric is a weak metric space that can be used to define both the Hilbert and the Thompson metrics, and is often studied in their context [22, 28] and in the context of flags of polytopes [7]. Because the Funk metric is non-symmetric, it induces a reverse metric called the reverse Funk metric. We contribute a proof that the minimum radius ball problem in these weak metrics are LP-type.
@@ 2.1 Metric Spaces
When all of the above properties are satisfied except for symmetry, the space is called a weak metric. Given a metric space, $(\mathbf{X},d)$, the closed ball of radius $r$ around a point $p$ generalizes the Euclidean idea of a circle to an arbitrary metric space. We define it here:
@@
$B(p,r)=\{q\in\mathbf{X}\mid d(p,q)\leq r\}$
@@ 2.2 The Hilbert and Thompson Metrics, and the Funk Weak Metric
The Hilbert and Thompson metrics, and the Funk weak metrics, are defined on the points in the interior of a convex body $\Omega$ in $\mathbb{R}^{d}$, where $\Omega$ is a closed and bounded convex set. Unless otherwise stated, we assume that $\Omega$ is a convex polygon with $m$ edges. We let $\partial\kern 1.0pt\Omega$ refer to the boundary of $\Omega$. Given any pair of points $p$ and $q$ within the interior of $\Omega$, we define $\overline{pq}$ to be the cord of $\Omega$ through those two points, and $\chi(p,q)$ to be the directed ray from $p$ through $q$. Unless otherwise stated, when we take $p,q\in\Omega$, we mean the interior of $\Omega$.
@@
Given two points $p,q$ in a convex polygon $\Omega$ in $\mathbb{R}^{n}$ such that $\chi(p,q)$ intersects $\partial\kern 1.0pt\Omega$ at a point $q^{\prime}$ on $\partial\kern 1.0pt\Omega$, we define the Funk weak metric to be:
@@
Given two distinct points $p,q$ in a convex polygon $\Omega$ in $\mathbb{R}^{n}$, let $p^{\prime}$ and $q^{\prime}$ denote the endpoints of the chord $\overline{pq}$ on $\partial\kern 1.0pt(\Omega)$, so that the points lie in order $\langle p^{\prime},p,q,q^{\prime}\rangle$, the Hilbert distance between $p$ and $q$, $H_{\Omega}(p,q)$, is:
@@
Since the product in the definition of the Hilbert metric is the cross ratio, which is preserved under projective transformations, it follows that the Hilbert metric is invariant under projective transformations[19]. It is also worth noting that straight lines are geodesics in the Hilbert metric, though not all geodesics are straight lines, and that the Hilbert metric satisfies all the properties of a metric, including symmetry and the triangle inequality. Nielsen and Shao showed how to algorithmically compute Hilbert balls with the aid of spokes.
Definition 2.5 (Spoke).
Given $p\in\Omega$ a spoke through $p$ from a vertex $v$ of $\Omega$ is $\overline{pv}$.
Succinctly, a Hilbert Ball of radius $r$ around a point $p$ can be constructed by computing all the points that are a distance of $r$ away from $p$ along the spokes of $p$ from the vertices of $\Omega$ and forming their convex hull (see Figure 1(c)).
@@
Hilbert balls have at most $O(m)$ sides.
@@
Given a bounded closed convex body $\Omega$ in $\mathbb{R}^{n}$ and two points $p,q\in\operatorname{int}(\Omega)$, let $p^{\prime}$ and $q^{\prime}$ denote the endpoints of the chord $\overline{pq}$ on $\partial\kern 1.0pt(\Omega)$, so that the points lie in order $\langle p^{\prime},p,q,q^{\prime}\rangle$, the Thompson distance between $p$ and $q$ $T_{\Omega}(p,q)$ is:
$T_{\Omega}(p,q)\leavevmode\nobreak\ =\leavevmode\nobreak\ max\left(\ln\frac{\|q-p^{\prime}\|}{\|p-p^{\prime}\|},\ln\frac{\|p-q^{\prime}\|}{\|q-q^{\prime}\|}\right),$
@@
Balls in the Thompson metric are convex polygons with $O(m)$ sides.
Proof 2.9.
This follows directly from the fact that balls in the Funk metric are homotheties (the reverse is flipped) of $\Omega$ with $O(m)$ sides [22] (see Figure 1 (a) and (d)) and Thompson balls are their intersections (see Figure 1(b)).
Theorem 2.10.
The topology induced by the Hilbert or Thompson metrics, as well as the Funk weak metrics in a bounded convex domain $\Omega\subset\mathbb{R}^{n}$ coincides with the Euclidean topology in that domain.
Proof 2.11.
This is a well known result in the field. For reference, this follows from Theorem $2.1$ in [3] where the authors show that the Hilbert and Euclidean metrics induce the same topology on bounded convex domains in $\mathbb{R}^{n}$. In addition, refer to Proposition $6.1$ in "From Funk to Hilbert Geometry" for Funk weak metrics.
From this we immediately have the following corollary.
Corollary 2.12.
Any Hilbert or Thompson metric space, as well as any Funk weak metric space, has the Heine-Borel property.
@@ 3 Minimum Enclosing Radius Balls and the Heine-Borel Property
Figure 2: This is $\mathbb{R}^{2}$ with several closed balls removed. There is no Euclidean minimum radius ball through $x,y,z$.
@@
For an example of a metric space without the minimum ball property see Figure 2. In this situation several closed disks were removed from $\mathbb{R}^{2}$ with the Euclidean distance.
@@
$S:=\{s\in\mathbb{R}^{\geq 0}\mid\text{ there exists a closed ball of radius }s\text{ enclosing }H\}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ r:=\inf S$
Note that $S$ is bounded below by 0 and is nonempty because $H$ is bounded as a consequence of being finite. So the infimum is defined and finite. Thus, to show there exists a ball of minimal radius, we want to show that there exists a ball of radius $r$ enclosing $H$. Define the set of centers $C_{n}:=\{p\in\mathbf{X}:H\subset B(p,r+1/n)\}$. For all $n\in\mathbb{N}$ we have $C_{n}\supset C_{n+1}$ because for all $p\in C_{n+1}$ we have $H\subset B(p,r+1/(n+1))\subset B(p,r+1/n)$ and thus $p\in C_{n}$ as well.
Each $C_{n}$ is bounded since for any $h\in H$ and any $p\in C_{n}$, $d(p,h)\leq r+1/n$.
We will now show that each $C_{n}$ is closed. Let $h\in H$ and $\{p_{i}\}_{i\in\mathbb{N}}\subset C_{n}$ such that $p_{i}\rightarrow p$, $p\in\mathbf{X}$. Then for any $\varepsilon>0$, there exists $i$ such that $d(p,p_{i})<\varepsilon$. So by the triangle inequality, $d(p,h)\leq d(p,p_{i})+d(p_{i},h)\leq(r+1/n)+\varepsilon$ for any $\varepsilon$, so $d(p,h)\leq r+1/n$. So $h\in B(p,r+1/n)$ for any $h\in H$. Thus $p\in C_{n}$. So $C_{n}$ contains all of its limit points and is thus closed.
Define $D:=\bigcap_{n}C_{n}$. Since we are assuming the Heine-Borel property and showed each $C_{n}$ is closed and bounded, each $C_{n}$ is compact. Because $\{C_{n}\}$ is a decreasing nested sequence of compact sets of $\mathbf{X}$ with the metric topology, by Cantor’s intersection theorem $D$ is nonempty [25].
@@
Note that the converse does not hold. As a counterexample, any infinite set equipped with the discrete metric does not satisfy the Heine-Borel property, yet it satisfies the minimum ball property. Additionally, note that in the above proof of Theorem 3.3 we did not use the symmetric property of the metric space, only the triangle inequality. As such, given the same definition of a ball Definition 2.2 and Definition 3.1 e.g. fixing the direction of the ball, we contribute the following corollary.
@@
We now introduce the characterization of an LP-type problem by Sharir and Welzl [26].
@@
Let $(\mathbf{X},d)$ be a metric space satisfying the minimum ball property. Let $\mathbf{X}$ have some well-ordering $\preceq$. Take a finite set $H\subset\mathbf{X}$. We define:
@@
Here, $f(G)=(\hat{r},\hat{p})$ gives us a unique minimum radius enclosing ball of $G$, by taking the minimum center with respect to the well-ordering. We call this ball $B_{G}=B(\hat{p},\hat{r})$. Note that if $f(G)=(r,p)$, there may exist other points $p^{\prime}$ such that $G\subset B(r,p^{\prime})$ as well; however, $f$ defines only one minimum radius enclosing ball. Note also that if $\mathbf{X}=\mathbb{R}^{n}$ then by Proposition 3.10 we can take the lexicographic usual total order and $f$ will still be well-defined.
@@
The problem of finding the minimum radius ball of a set of points in a metric space that satisfies the minimum ball property is LP-type.
@@
To see that $f$ satisfies locality, consider a finite set $H\subset\mathbf{X}$. Take $F,G\subset H$ and $x\in H$ such that $F\subset G\subset H$ and $f(F)=f(G)=f(F\cup\{x\})$. It follows that $B_{F}=B_{G}=B_{F\cup\{x\}}$. Hence, $x\in B_{G}$ and $F\subset G\cup\{x\}\subset B_{G}=B_{F}$. Thus, by monotonicity, $f(F)\leq f(G\cup\{x\})\leq f(G)=f(F)$, and we conclude that $f(F)=f(G\cup\{x\})$.
@@
Finding the minimum radius ball of a set of points in a weak metric space that satisfies the minimum ball property is LP-type.
Note that a well-ordering is not always necessary and can be replaced with a total ordering, such as a lexicographic ordering.
@@
If $X=\mathbb{R}^{k}$ for $n<\infty$ then the well-order $\preceq$ on $\mathbf{X}$ can be replaced with the lexicographical usual total order $\leq$ from $\mathbb{R}^{k}$.
@@
Recall, the following notation and results from Theorem 3.3.
$C_{n}:=\{p\in\mathbf{X}:H\subset B(p,r+1/n)\}\qquad\text{and}\qquad D:=\bigcap_{n}C_{n}$
We saw that each $C_{n}$ is closed and bounded with respect to the metric topology. Thus it is compact in the standard (metric) topology in $\mathbb{R}^{k}$, and thus $D$, the set of centers of minimum radius enclosing balls is also nonempty and compact in this topology.
@@
Corollary 3.12.
The minimum radius ball problem in the Hilbert and Thompson metrics, as well as the Funk weak metrics, is LP-type.
Proof 3.13.
Note that Corollary 2.12, Proposition 3.3, and Theorem 3.7 together imply this result for the Hilbert and Thompson metrics. Note secondly that Corollary 10, Corollary 3.5, and Corollary 3.9 together imply the result for the Funk weak metrics.